OpenStudy (anonymous):

What is the volume of a cube with sides that are 9 m long? A. 27 cubic meters B. 81 cubic meters C. 486 cubic meters D. 729 cubic meters

jimthompson5910 (jim_thompson5910):

Volume of Cube = (side length)^3 which is the same as saying Volume of Cube = (side length)*(side length)*(side length)

OpenStudy (anonymous):

the answer i got was 531441

jimthompson5910 (jim_thompson5910):

that's way too big

OpenStudy (whpalmer4):

\[V = s^3 = s*s*s\]\[s=9\]\[V = 9*9*9 = \]Units will be \(m*m*m = m^3\)

OpenStudy (anonymous):

i got 6561

OpenStudy (whpalmer4):

How are you multiplying? Calculator, pencil & paper, abacus, gathering many friends to borrow fingers and toes? :-)

OpenStudy (whpalmer4):

What is 9*9?

OpenStudy (anonymous):

i'm using calculater and 9*9 is 81

OpenStudy (whpalmer4):

okay, what is 81*9?

OpenStudy (anonymous):

729

OpenStudy (whpalmer4):

Okay, so \[V = s^3 = 9^3 = 9*9*9 = 729\]

OpenStudy (anonymous):

Length (9m) x Width (9m) x Height (9m) = Area(Volume) of Cube at approximately 729m³

OpenStudy (anonymous):

oh ok thanks i was adding 9 6 times

OpenStudy (whpalmer4):

@butterfly5 there's no "approximately" here, it is exactly 729 cubic meters

OpenStudy (anonymous):

lol